Reece earned $250 000 in 2014 as a celebrity endorser and spokesperson. She spent 85 percent of her earnings. In 2015, with a new contract, she earned $750 000 per year. Her consumption increased by $250 000. Which of the following statements about her propensity to consume is NOT accurate?

a. Her 2015 marginal propensity to consume is 0.50.

b. Her 2014 average propensity to consume is 0.85.

c. Her average propensity to consume fell between 2014 and 2015.

d. Her marginal propensity to save falls to 0.15 in 2015. e. None of the above

Which of the following describes marginal propensity to consume (MPC) and average propensity to consume (APC)?

a. MPC is less than 1, but APC can be either greater or less than 1. b. Both MPC and APC are always less than 1.
c. MPC is less than 1, but APC is always equal to 1.

d. MPC and APC are always greater than or equal to 1.

e. None of the above.

  1. Which of the following changes in disposable income would lead to the greatest increase in consumption?

    a. a $40 000 increase in disposable income, if MPC equals 0.5

    b. a $30 000 increase in disposable income, if MPC equals 0.6

    c. a $24 000 increase in disposable income, if MPC equals 0.75

    d. a $60 000 increase in disposable income, if MPC equals 0.25 e. None of the above
Answer #1

Q 1) The following table summarizes the consumption and income in 2014 and 2015:

INCOME CONSUMPTION
2014 250000 212500
2015 750000

462500

Marginal propensity to consume = change in consumption / change in income

So, marginal propensity to consume in 2015 = (462500 - 212500) / (750000 - 250000)

Marginal propensity to consume in 2015 = 250000 / 500000 = 1/2 = 0.5

Marginal propensity to consume + Marginal propensity to save = 1

So, 0.5 + Marginal propensity to save = 1

Marginal propensity to save = 1 - 0.5 = 0.5

So, statement d is not accurate that is marginal propensity to save falls to 0.15 in 2015. Marginal propensity to save is 0.5 in 2015.